# Writing – review & editing, Writing – original draft, Visualization, Data curation ,Investigation, Formal analysis, Validation, Software, Methodology, Conceptualization

### v2

@@ 5.4 Attack Quality
In highly constrained datasets like the ICU dataset where many features cannot be modified and over 30% of the features are categorical, transferability attacks tend to produce perturbations of similar magnitude to query-based attacks on categorical features, making query-based approaches competitive in these scenarios. The limited number of editable features forces both query- and transferability-based attacks to make fewer changes, causing less anomalous behavior overall. The random-based selection attack is an exception, as it compromises coherence by arbitrarily modifying features, often resulting in anomalous values and noticeable changes in SHAP values.
These findings suggest that defense mechanisms should consider not only the attack strategy but also the specific dataset and model vulnerabilities.
Our evaluation of attack quality—employing both the anomaly detection rate to assess sample coherence and the importance-based anomaly detection rate to evaluate the impact of adversarial samples on the target model’s decision-making process—reveals complementary insights. The Importance-Based Anomaly Detection Rate (computed based on SHAP values, see Table 4), can help identify adversarial samples that appear normal when evaluating just their feature values. For example, on the Hate dataset, the HopSkipJump attack only produced anomalous values when performed against the GB and LGB target models (see Fig. 3). However, when evaluating anomalies in SHAP values, our proposed metric captured anomalies when attacking the XGB and RF target models as well (see Table 4. On the VideoTQ dataset, we observed the opposite phenomenon; here, traditional anomaly detection (specifically using the AE model) captured anomalies for all of the transferabilty-based attacks (see Fig. 3), while the SHAP values were not anomalous in many cases (see Table 4).
@@ 5.5 Effectiveness of the Class-Specific Anomaly Detection (CSAD) Approach
Table 5 presents the results of two approaches for computing the anomaly detection rate: (1) the CSAD approach, which measures the anomalousness of SHAP values in adversarial samples in relation to benign SHAP values of the same predicted class only, and (2) the standard approach, which compares adversarial samples to the entire benign SHAP value distribution across all classes, enabling a comparative analysis of detection effectiveness between class-specific and general anomaly detection methods.
@@ 6 Discussion
Regarding anomaly detection models, the neural network-based AE detector identified more adversarial samples as anomalous than the tree-based IF alternative (as expected). These empirical results validate that AEs better capture the complex patterns and relationships in the data, making them more sensitive to subtle adversarial manipulations in tabular datasets.
@@ 7 Conclusion
The results obtained using diverse target models and datasets highlight distinct trade-offs: query-based attacks, while highly effective, often sacrifice coherence by causing significant distortions, and require substantial computational resources. Transferability-based attacks, though less successful in some cases, achieve greater coherence and consistency with fewer feature modifications, resulting in subtler impact on model decision-making. This is evidenced by lower anomaly detection rates and minimal changes in feature importance.
@@
Future work could investigate adversarial robustness in more complex tabular ML pipelines, including scenarios where models incorporate built-in defenses against such attacks or where attackers does not have full knowledge of the target model’s features. By building on the insights and methodologies presented here, researchers can further refine both attack and defense strategies in the tabular domain.
@@ Appendix A Pseudocode for Tabular Attacks
In this research, we applied two types of query-based attacks: the Boundary and HopSkipJump attacks, along with five versions of transferability-based attack, all adapted for the tabular domain and outlined below.
In the Boundary attack presented in Algorithm 1, the attacker randomly selects a sample from the defined target class, regardless of its proximity to the original sample (lines 2-3). The attacker then optimizes the modified sample’s values using an orthogonal perturbation, which iteratively adjusts the feature values to minimize the model’s confidence in the original class while ensuring that the predicted class is the target class (lines 4-6). During the crafting process, tabular constraints are applied to the sample (see Algorithm 2), projecting each feature value onto its closest legitimate value to ensure input validity (line 7).
Input: The target model $M$, original sample $x$, original sample’s true label $y$, set of immutable features $I$, set of each feature constraints $Constraints$, and the similarity threshold $\alpha$
[…]
2 while $i<max\_init\_steps\land M(x^{adv})=y$ do
[…]
end while
[…]
      5while $j<max\_similarity\_steps$ do
[…]
       end while
[…]
end while
8return $x^{adv}$
Algorithm 1 The Boundary adversarial attack for tabular data
Input: The original sample $x$, adversarial sample $x^{adv}$, set of immutable features $I$, and the set of each feature constraints $Constraints$
1 Procedure Tabular_Modify(x, $x^{adv}$, I, Constraints, regression_models)
@@ 5 Correct dependent features (regression_models)
Algorithm 2 A generic procedure for applying constraints of tabular data domains
In contrast, in the HopSkipJump attack, which is presented in Algorithm 3, the attacker employs binary search and proceeds in the target class gradients’ direction to find a sample from the defined target class with the closest proximity to the original sample (lines 5-10). During the binary search, tabular constraints are applied to the sample (see Algorithm 2), projecting each feature value onto its closest legitimate value to ensure input validity (line 11).
Input: The target model $M$, original sample $x$, original sample’s true label $y$, set of immutable features $I$, set of each feature constraints $Constraints$, and the similarity threshold $\alpha$
@@
2 while $i<max\_init\_steps\land M(x^{adv})=y$ do
@@ 4 $x^{adv}\leftarrow$ Binary_Search(x, $x^{adv})$
      6while $j<max\_similarity\_steps$ do
[…]
            8 $update\leftarrow$ Compute_Update(x, $x^{adv}$)
            9 $potential\_sample\leftarrow$ run step size search ($x^{adv},update$) until predicts satisfying
            10 $x^{adv}\leftarrow potential\_sample$
@@
12return $x^{adv}$
@@
In our evaluation, we implemented untargeted transferability-based gradient attacks based on the architecture in [36, 21], and the pseudocode is presented in Algorithm 4. The attacker first selects the feature likely to have the most impact on the model’s prediction, excluding immutable features, i.e., features that cannot be changed by the attacker (line 4). The selection, which is presented in lines 9-13, can be made by: 1) selecting $k$ random features and the $n$ features that are most correlated with them  [36]. 2-4) using a feature importance technique to select the $k$ most important features [21] and the $n$ features that are most correlated with them. The feature importance technique can be performed by training an ML classifier (such as GB, XGB, LGB, and RF models) and extracting what they consider the most important features [21]. After selecting features to perturb, a perturbation vector $p$ is calculated for the selected features by using an optimizer that minimizes the adversary’s objective function (line 5). Finally, tabular constraints are applied to the sample (see Algorithm 2), projecting each feature value onto its closest legitimate value to ensure input validity (line 7).
Input: The target model $M$, original sample $x$, original sample’s true label $y$, adversary’s objective loss function $L^{adv}$, set of immutable features $I$, set of each feature constraints $Constraints$, maximum allowed $L_{0}$-perturbation noise $\lambda$, number of most important features to perturb $k$, and number of features most correlated with the $k$ features chosen $n$
1 $x^{adv}\leftarrow x$
@@ 2 $F\leftarrow\emptyset$
3 while $M^{\prime}(x^{adv})=y\land|F|<\lambda$ do
@@
      5 $p\leftarrow$ Compute_Perturbation($M^{\prime},L^{adv},x^{adv},F$)
      6 $x^{adv}_{i}\leftarrow x^{adv}_{i}+p$
      7 $x^{adv}\leftarrow$ Tabular_Modify($x,x^{adv},I,Constraints$)
8return $x^{adv}$
Procedure Select_Features($M^{\prime},x^{adv},I,F,k,n$)
      9 •Random-Based Selection
      10 •GB Feature Importance-Based Selection
      11 •XGB Feature Importance-Based Selection
      12 •LGB Feature Importance-Based Selection
      13 •RF Feature Importance-Based Selection

### v3

@@ 5.4 Attack Quality
In highly constrained datasets like the ICU dataset where many features cannot be modified and over 30% of the features are categorical, transferability attacks tend to produce perturbations of similar magnitude to query-based attacks on categorical features, making query-based approaches competitive in these scenarios. The limited number of editable features forces both query- and transferability-based attacks to make fewer changes, causing less anomalous behavior overall. The random-based selection attack is an exception, as it compromises coherence by arbitrarily modifying features, often resulting in anomalous values and noticeable changes in SHAP values. These findings suggest that defense mechanisms should consider not only the attack strategy but also the specific dataset and model vulnerabilities.
Our evaluation of attack quality—employing both the anomaly detection rate to assess sample coherence and the importance-based anomaly detection rate to evaluate the impact of adversarial samples on the target model’s decision-making process—reveals complementary insights. The Importance-Based Anomaly Detection Rate (computed based on SHAP values, see Table˜5), can help identify adversarial samples that appear normal when evaluating just their feature space values. For example, on the Hate dataset under the boundary attack against the XGB model, traditional anomaly detection (particularly with IF) achieved nearly perfect detection, whereas the SHAP-based anomaly rate remained very low (see Figs.˜3 and 5). In contrast, on the VideoTQ dataset under the LGB importance-based selection attack, the SHAP-based anomaly rate exceeded $90\%$, while traditional anomaly detection failed almost entirely (see Figs.˜3 and 5).
@@ 5.5 Effectiveness of the Class-Specific Anomaly Detection (CSAD) Approach
Table˜6 presents the results of two approaches for computing the anomaly detection rate: (1) the CSAD approach, which measures the anomalousness of SHAP values in adversarial samples in relation to benign SHAP values of the same predicted class only, and (2) the standard approach, which compares adversarial samples to the entire benign SHAP value distribution across all classes, enabling a comparative analysis of detection effectiveness between class-specific and general anomaly detection methods.
@@
Statistical validation using McNemar exact tests (with Holm correction) on the importance-based anomaly rate confirmed that CSAD consistently and significantly outperforms the standard approach across all datasets ($p<0.001$). Effect sizes were uniformly large (Cohen’s $g\approx 1.0$), indicating a strong and systematic advantage of CSAD. Complementary analyses of average anomalous SHAP features per sample (Wilcoxon signed-rank, Holm-corrected) led to the same conclusion, with rank-biserial correlations also close to 1.0. These results highlight that the superiority of CSAD is statistically significant and practically substantial, underscoring its robustness across datasets and attack scenarios. Full test results are provided in Appendix C.
@@ 6 Discussion
Regarding anomaly detection models, IF consistently outperformed AE across datasets, particularly under boundary and random-based attacks. While AEs are designed to capture complex feature interactions, in our evaluation their effectiveness was limited and inconsistent, with meaningful detection primarily observed under the query-based scenarios. These results suggest that tree-based methods like IF may provide stronger baselines for tabular anomaly detection, while neural approaches such as AEs require further refinement to reliably capture adversarial perturbations.
The generalizability of our methodological contributions extends naturally to multi-class classification scenarios, despite our empirical evaluation focusing on binary classification datasets. Our regression-based perturbation technique preserves feature relationships regardless of the number of target classes, as it maintains coherence based on feature dependencies rather than classification boundaries. The CSAD approach inherently supports multi-class settings by comparing adversarial samples against benign samples from the same predicted class, regardless of whether they are binary or multi-class. Similarly, the SHAP-based evaluation metrics apply directly to multi-class models, as SHAP provides feature importance scores for each class prediction. The focus on binary classification enables thorough analysis of the proposed evaluation framework while maintaining computational feasibility for comprehensive comparison across multiple attack strategies and models; the use of established benchmark datasets facilitates direct comparison with prior tabular adversarial research.
Finally, we note an important limitation of our regression-based perturbation technique. In our evaluation, the method was applied to datasets where at least some feature dependencies could be identified in advance, either from domain knowledge or through direct analysis of the training data. In scenarios where such dependencies are mutual, non-trivial, or entirely latent, an additional preliminary stage is required to infer them before perturbation. This detection phase could employ statistical dependency measures, feature clustering, or model-based approaches to uncover hidden relationships. While our framework is theoretically compatible with integrating such a stage, its design and implementation remain an open avenue for future research.
@@ 7 Conclusion
The results obtained using diverse target models and datasets highlight distinct trade-offs: query-based attacks, while highly effective, often sacrifice coherence by causing significant distortions, and require substantial computational resources. Transferability-based attacks, although less successful in some cases, achieve greater coherence and consistency with fewer feature modifications, resulting in subtler impact on model decision-making. This is evidenced by lower anomaly detection rates and minimal changes in feature importance.
@@
Future work could perform ablation studies on alternative preprocessing pipelines to better understand their impact on adversarial attack effectiveness and model robustness. Another promising direction is to extend the regression-based perturbation framework to scenarios where feature dependencies are mutual, non-trivial, or entirely latent. This would require incorporating a preliminary dependency discovery stage, using, for example, statistical measures, feature clustering, or model-based inference, to identify and model hidden relationships before perturbation. Investigating adversarial robustness in such settings, including cases where models incorporate built-in defenses or where attackers lack full knowledge of the target model’s features, could provide deeper insights and further refine both attack and defense strategies in the tabular domain.
8 Data and Code Availability
To promote reproducibility and facilitate future research, we provide open access to all research artifacts generated in this study. This includes: (1) the preprocessed benchmark datasets, (2) the implementation code for the evaluated attack methods and defense mechanisms, (3) the trained target models, and (4) the adversarial samples generated. All resources are available at https://github.com/yael87/Framework-for-adversarial-in-tabular-domain.
@@ Appendix A Pseudocode for Tabular Attacks
In this research, we applied two types of query-based attacks: the boundary and HopSkipJump attacks, along with five versions of transferability-based attacks, all adapted for the tabular domain and outlined below. All attacks include a generic constraint modification procedure to ensure that the generated adversarial samples remain valid in the tabular data domain.
A.1 Tabular Constraint Modification
The tabular domain requires specific constraint handling to ensure that adversarial samples maintain realistic and valid feature values. Our generic constraint modification procedure Algorithm˜1 addresses four key aspects: (1) clipping feature values to their valid ranges (e.g., age cannot be negative), (2) imposing data type constraints (e.g., integer values for discrete features), (3) preserving immutable features that cannot be modified by an attacker (e.g., gender, age), and (4) correcting dependent features using pretrained regression models to maintain realistic relationships within the rest of the sample.
Input: The original sample $x$, adversarial sample $x_{adv}$, set of immutable features II, the set of each feature constraints ConstraintsConstraints, and set of regression models regression_modelsregression\_models(if exist)
1 Procedure Tabular_Modify(x, $x_{adv}$, I, Constraints, regression_models)
@@ 5 Correct dependent features (regression_models)
Algorithm 1 A generic procedure for applying constraints in tabular data domains: clipping each feature value to its valid range, imposing realistic values according to feature constraints (e.g., integer, positive, normalized), ensuring immutable features maintain their original values, and correcting dependent features based on corresponding regression models.
A.2 The Boundary Attack
In the boundary attack presented in Algorithm˜2, the attacker randomly selects a sample from the defined target class, regardless of its proximity to the original sample (lines 2-3). The attacker then optimizes the modified sample’s values using an orthogonal perturbation, which iteratively adjusts the feature values to minimize the model’s confidence in the original class while ensuring that the predicted class is the target class (lines 4-6). During the crafting process, tabular constraints are applied to the sample (see Algorithm˜1), projecting each feature value onto its closest legitimate value to ensure input validity (line 7).
Input: The target model $M$, original sample $x$, original sample’s true label $y$, set of immutable features II, set of each feature constraints ConstraintsConstraints, and the similarity threshold $\alpha$
@@ A.2 The Boundary Attack
2 while $i<max\_init\_steps\land M(x_{adv})=y$ do
@@
4while $i<max\_iteration\_steps\land M(x_{adv})=y\land$ Similarity($x,x_{adv})<\alpha$ do
 5while j<max_similarity_stepsj<max\_similarity\_steps do
@@
8return $x_{adv}$
Algorithm 2 The boundary adversarial attack for tabular data
A.3 The HopSkipJump Attack
In the HopSkipJump attack, which is presented in Algorithm˜3, the attacker employs binary search and proceeds in the target class gradients’ direction to find a sample from the defined target class with the closest proximity to the original sample (lines 5-10). During the binary search, tabular constraints are applied to the sample (see Algorithm˜1), projecting each feature value onto its closest legitimate value to ensure input validity (line 11).
Input: The target model $M$, original sample $x$, original sample’s true label $y$, set of immutable features II, set of each feature constraints ConstraintsConstraints, and the similarity threshold $\alpha$
[…]
2 while $i<max\_init\_steps\land M(x_{adv})=y$ do
[…]
 4 $x_{adv}\leftarrow$ Binary_Search(x, $x_{adv})$
end while
[…]
 6while j<max_similarity_stepsj<max\_similarity\_steps do
[…]
    8 $update\leftarrow$ Compute_Update(x, $x_{adv}$)
    9 $potential\_sample\leftarrow$ run step size search ($x_{adv},update$) until predicts satisfying
    10 $x_{adv}\leftarrow potential\_sample$
   end while
[…]
end while
12return $x_{adv}$
@@ 10 $x_{adv}\leftarrow potential\_sample$
A.4 The Transferability-Based Attacks
The transferability-based attacks leverage surrogate models to generate adversarial samples that transfer to black-box target models. These attacks iteratively select and perturb features based on different selection strategies until successful misclassification is achieved or the maximum perturbation budget is exhausted.
In our evaluation, we implemented untargeted transferability-based gradient attacks based on the architecture in [8, 9], with the pseudocode presented in Algorithm˜4. In each iteration of these attacks, the attacker first selects features to pertrube, excluding immutable features that cannot be changed by the attacker (line 4). The feature selection process (lines 10-14) differs across attack variants and follows one of five strategies: 1) Random-based selection: selecting $k$ random features and the $n$ features that are most correlated with them. 2-5) Importance-based selection: selecting the $k$ features most likely to impact the model’s prediction and the $n$ features most correlated with them. Feature importance is determined using SHAP values computed from trained ML classifiers, with each attack variant using a different classifier (gradient boosting, XGBoost, LightGBM, or random forest) to compute these SHAP values  [9].
After selecting features to perturb, a perturbation vector $p$ is calculated for the selected features (line 5) by using an optimizer that minimizes the adversary’s objective function $L_{adv}$ which is defined in Section˜4 of the paper. Finally, tabular constraints are applied to the sample using Algorithm˜1, to ensure that the adversarial sample maintains valid feature values (line 7).
The feature selection process is detailed in the Select_Features procedure, which maintains a set $F$ of already-selected features to avoid redundant selections across iterations.
Input: The surrogate model $M^{\prime}$, original sample $x$, original sample’s true label $y$, adversary’s objective loss function $L_{adv}$, set of immutable features II, set of constraints for each feature $Constraints$, maximum allowed $L_{0}$-perturbation noise $\lambda$, number of most important features to perturb $k$, and number of features most correlated with the $k$ features chosen $n$
1 $x_{adv}\leftarrow x$
@@ 2 $F\leftarrow\emptyset$
3 while $M^{\prime}(x_{adv})=y\land|F|<\lambda$ do
@@
 5 $p\leftarrow$ Compute_Perturbation($M^{\prime},L_{adv},x_{adv},F$)
 6 $x_{adv}\leftarrow x_{adv}+p$
 7 $x_{adv}\leftarrow$ Tabular_Modify($x,x_{adv},I,Constraints$)
8return $x_{adv}$
Procedure Select_Features($M^{\prime},x_{adv},I,F,k,n$)
 9 Input: Surrogate model $M^{\prime}$, current adversarial sample $x_{adv}$, immutable features II, previously selected features $F$, number of features to select $k$, number of correlated features per selection $n$
 10 1) Random-Based Selection: Select $k$ random features from available features not in $F\cup I$, then select $n$ most correlated features for each
 11 2) GB Feature Importance-Based Selection: Compute SHAP values using gradient boosting model, select top $k$ features not in $F\cup I$, then $n$ correlated features
 12 3) XGB Feature Importance-Based Selection: Compute SHAP values using XGBoost model, select top $k$ features not in $F\cup I$, then $n$ correlated features
 13 4) LGB Feature Importance-Based Selection: Compute SHAP values using LightGBM model, select top $k$ features not in $F\cup I$, then $n$ correlated features
 14 5) RF Feature Importance-Based Selection: Compute SHAP values using random forest model, select top $k$ features not in $F\cup I$, then $n$ correlated features
 15 return Set of newly selected features